OpenStudy (anonymous):

write the following answers in ordered pairs (x,y) 3.Find the midpoint between the points (3, 1) and (1, 5). 4.Find the midpoint between the points (5, 2) and (-1, -2). please show steps

OpenStudy (anonymous):

mid point m : mx = (x1+x2)/2 my = (y1 +y2)/2 so 3. mx = (3+1)/2 = 2 my = (1+5)/2 = 3 (2,3) 4. mx = (5-1)/2 = 2 my = (2-2)/2 = 0 (2,0)

OpenStudy (anonymous):

the mid point formula is ((x1+x2)/2, (y1+y2)/2)

OpenStudy (anonymous):

hello